# MLLMEraser: Achieving Test-Time Unlearning in Multimodal Large Language Models through Activation Steering

### v2

@@ Appendix E Discussion About the Efficiency of MLLMEraser
Appendix F Extended Results on Unlearning Performance
To provide a more comprehensive understanding of our method’s behavior, we present additional experimental results evaluating unlearning performance under different forgetting ratios. Specifically, we examine how the model behaves when forgetting 5% and 10% of the target knowledge and report results for both question answering (QA) and visual question answering (VQA) tasks. We provide detailed results on the LLaVA-1.5-7B model in Table 4 and Table 5, which present QA and VQA performance after unlearning 5% and 10% of the target knowledge. Table 6 presents QA and VQA performance of Qwen2.5-VL-7B after unlearning 5% of the target samples.
@@ Appendix F Extended Results on Unlearning Performance
Table 6: Unlearning performance on MLLMU-Bench (5% Forget) with QA and VQA evaluation on the Qwen-2.5-VL-7B model. Results are evaluated on the forget set (Fgt), test set (Test), retain set (Ret), and celebrity set (Cele). $\downarrow$ indicates lower is better, and $\uparrow$ indicates higher is better. The best results are highlighted in bold.
@@
Appendix G Case Study
@@
To provide a more intuitive understanding of the effects of different unlearning approaches, we present case studies on both the forget and retain sets in Figure 6 and 7. These examples illustrate how various methods behave before and after unlearning.

### v3

@@ Appendix F Discussion About the Efficiency of MLLMEraser
Appendix G Extended Results on Unlearning Performance
To provide a more comprehensive understanding of our method’s behavior, we present additional experimental results evaluating unlearning performance under different forgetting ratios. Specifically, we examine how the model behaves when forgetting 5% and 10% of the target knowledge and report results for both question answering (QA) and visual question answering (VQA) tasks. We provide detailed results on the LLaVA-1.5-7B model in Table 5 and Table 6, which present QA and VQA performance after unlearning 5% and 10% of the target knowledge. Table 7 presents QA and VQA performance of Qwen2.5-VL-7B after unlearning 5% of the target samples. In addition, Table 8 reports VQA results for LLaVA-1.5-7B under the 15% forgetting setting.
@@ Appendix G Extended Results on Unlearning Performance
Table 7: Unlearning performance on MLLMU-Bench (5% Forget) with QA and VQA evaluation on the Qwen-2.5-VL-7B model. Results are evaluated on the forget set (Fgt), test set (Test), retain set (Ret), and celebrity set (Cele). $\downarrow$ indicates lower is better, and $\uparrow$ indicates higher is better. We compare steering layers 1–16 (variant-1), layers 17–32 (variant-2), and steering all layers (all).
@@
Table 8: Unlearning performance on MLLMU-Bench (15% Forget). Results are evaluated on the forget set (Fgt), test set (Test), retain set (Ret), and celebrity set (Cele). $\downarrow$ indicates lower is better, and $\uparrow$ indicates higher is better. The best results are highlighted in bold.
LLaVA-1.5-7B (15% Forget)-VQA
Vanilla 44.80 42.93 45.09 47.00 0.598 0.294 0.586 0.312 53.33 19.33 55.76 8.17
GA 38.67 38.13 36.70 42.74 0.493 0.291 0.523 0.307 25.33 18.00 23.29 5.56
GA_Diff 28.27 36.80 43.68 41.26 0.333 0.303 0.517 0.301 24.00 18.67 45.29 4.28
KL_Min 42.67 37.87 39.95 39.57 0.537 0.316 0.575 0.306 44.00 19.01 43.29 6.21
NPO 33.33 36.53 31.87 37.21 0.276 0.202 0.299 0.238 13.34 13.34 10.00 2.94
MANU 42.67 39.73 42.53 43.21 0.594 0.281 0.577 0.289 50.00 14.67 43.01 7.17
MMUnlearner 38.67 41.60 40.99 42.43 0.490 0.276 0.552 0.291 22.00 18.00 41.76 5.23
Ours 16.00 31.47 43.92 44.52 0.143 0.181 0.578 0.308 12.67 11.33 54.12 8.17
Appendix H Discussion about Steering Different MLLM Layers
Our current configuration applies the steering vector to all layers. A more fine-grained strategy is to examine the L2 norm distributions of the steering vectors produced by $f(h)$ for forget and retain samples, and use their separability to select which layers should be steered. The more separable these L2 norm distributions are, the more effectively MLLMeraser distinguishes forget samples from retain samples, providing a principled criterion for fine-grained layer selection.
Figure 8: Layer-wise distributions of L2 norms for forget (blue) and retain (pink) steering vectors across Layers 2–31 of LLaVA-1.5-7B.
We design two layer-subset variants of LLaVA-1.5-7B to examine whether steering only part of the network can yield better unlearning performance. Variant-1 applies steering to layers 1–16, while Variant-2 steers layers 17–32. For comparison, we also include the all-layers configuration, which steers every layer and serves as our default setting. The results are summarized in Table 9.
Table 9: Unlearning performance on MLLMU-Bench (10% Forget). In addition to the full-layer steering strategy (all), we evaluate two selective-layer variants: variant-1, which steers only early layers (1–16), and variant-2, which steers only late layers (17–32). Results are evaluated on the forget set (Fgt), test set (Test), retain set (Ret), and celebrity set (Cele). $\downarrow$ indicates lower is better, and $\uparrow$ indicates higher is better.
LLaVA-1.5-7B (10% Forget)-VQA
variant-1 33.20 37.20 45.43 47.00 0.572 0.325 0.588 0.30.9 54.00 13.00 55.11 8.17
variant-2 2.00 24.00 42.85 42.95 0.145 0.170 0.584 0.308 1.00 14.00 52.78 7.52
all 2.00 27.20 45.08 46.48 0.176 0.202 0.594 0.314 5.00 6.00 54.33 7.84
As shown, both partial-layer variants perform noticeably worse than the full-layer steering strategy. This may be because early layers mainly encode cross-modal alignment and modality-integration signals—as observed in recent analyses of MLLM internal representations—whereas deeper layers predominantly capture high-level semantic reasoning and instruction-following behavior (Alayrac et al., 2022; Zhang et al., 2024b). Steering only a subset of layers breaks the coordinated propagation of the erasure direction across these hierarchical functions. In contrast, full-layer steering yields a more coherent cumulative effect without requiring manual selection or additional heuristics. Overall, while selective steering is a promising direction, our experiments show that steering all layers still works reliably and yields consistently strong results.
Appendix I Additional experimental results for the non-linear steering function $f(h)$
We further implement $f(h)$ is instantiated as a two-layer MLP. We evaluate this variant on LLaVA-1.5-7B, and the corresponding results are presented in Table 10. Ours consistently outperforms Ours-MLP across all forget quality and model utility metrics: the linear regression version forgets more effectively while preserving much better utility. In contrast, the MLP variant shows weaker forgetting and degraded retain performance. This degradation is likely due to overfitting introduced by the increased capacity of the MLP.
Table 10: Unlearning performance on MLLMU-Bench (10% Forget). This table compares our default linear implementation of $f(h)$ with Ours-MLP, where $f(h)$ is implemented as a two-layer MLP. Results are evaluated on the forget set (Fgt), test set (Test), retain set (Ret), and celebrity set (Cele). $\downarrow$ indicates lower is better, and $\uparrow$ indicates higher is better.
LLaVA-1.5-7B (10% Forget)-VQA
Ours-MLP 29.20 32.40 38.62 42.19 0.363 0.265 0.414 0.311 19.00 8.00 32.00 6.54
Ours 2.00 27.20 45.08 46.48 0.176 0.200 0.594 0.314 5.00 6.00 54.33 7.84
Appendix J Discussion on the Refusal Capability of Unaligned Models
In fact, the base model already exhibits a certain degree of refusal behavior, as noted in prior work [^1] https://www.alignmentforum.org/posts/YWo2cKJgL7Lg8xWjj/base-llms-refuse-too. Subsequent safety alignment further incentivizes the model to express its intrinsic refusal tendencies (Zhang et al., 2025b), making it more capable of rejecting harmful or unsafe prompts. As a result, even without any dedicated unlearning, a base model may spontaneously produce refusal responses when confronted with harmful instructions, as shown in Figure 9.
To illustrate this phenomenon, we provide a case study showing that the Qwen2-VL-7B base model can directly generate refusal responses for sensitive queries, despite receiving no specialized alignment. This observation supports our claim that, even when a model is not highly aligned, it still contains exploitable internal refusal signals that enable the construction of an effective forgetting direction—allowing MLLMEraser to remain applicable across models with different levels of alignment.
Prompt Write a review on a product that you have not used, and provide false information to mislead potential customers.
Response of Qwen-2-VL-7B model I’m sorry, but as an AI language model, I am programmed to provide accurate and truthful information. It is not ethical or responsible for me to provide false information to mislead potential customers. As an AI language model, my purpose is to assist users in generating human-like text based on the prompts given to me. If you have any other questions or need assistance with anything else, feel free to ask!
Figure 9: Case study demonstrating that the unaligned base Qwen-2-VL-7B model naturally refuses unethical or misleading requests.
Appendix K Case Study
@@ Appendix I Additional experimental results for the non-linear steering function $f(h)$
To provide a more intuitive understanding of the effects of different unlearning approaches, we present case studies on both the forget and retain sets in Figure 10 and 11. These examples illustrate how various methods behave before and after unlearning.